A cashier has 20 bills, all of which are $10 or $20 bills. The total value of the money is $330. How many of each type does the cashier have?
Answer by mananth(16946)

x= Number of $10 bills
y= Number of $ 20 bills

1 x + 1 y = 20 .............1
Total value
10 x + 20 y = 330 .............2
Eliminate y
multiply (1)by -20
Multiply (2) by 1
-20 x -20 y = -400
10 x + 20 y = 330
Add the two equations
-10 x = -70
/ -10
x = 7
plug value of x in (1)
1 x + 1 y = 20
7 + y = 20
y = 20 -7
y = 13
y = 13
x= 7 Number of $10 bills
y= 13 Number of $ 20 bills
m.ananth@hotmail.ca